OpenStudy (anonymous):

You are baking cookies. The recipe calls for 3/4 cup of flour. You want to make 1/2 of a batch. How much flour do you need?

OpenStudy (anonymous):

A full batch requires 3/4 cup. But you want to make 1/2 of a batch. The key here is the "1/2 of". Whenever you see wording like this, multiplication is required. So you need to solve\[\frac{ 1 }{ 2 }\times \frac{ 3 }{ 4 } = ?\] Do you know how to multiply fractions?

OpenStudy (anonymous):

Great. What answer do you get?

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

Excellent. That's your answer. Well done.